OpenStudy (anonymous):

|(n+1)^(1/2) - n^(1/2)| = 1/((n+1)^(1/2) + n^(1/2)) Could someone explain this equality to me, having trouble. Thanks.

OpenStudy (anonymous):

\[ \sqrt{n+1}-\sqrt{n}=\frac{1}{\sqrt{n}+\sqrt{n+1}} \] No need for absolute value is the quantity is positive. Cross multuply and you get n+1 -n = 1 You are done

OpenStudy (anonymous):

I was trying to work out how you produce the RHS having been given the LHS

OpenStudy (anonymous):

Them tutors always like to make life difficult!

OpenStudy (anonymous):

\[ \sqrt{n+1}-\sqrt{n}=\frac{\left(\sqrt{n+1}-\sqrt{n}\right ) \left(\sqrt{n+1}+\sqrt{n}\right)}{\sqrt{n+1}+\sqrt{n}}= \frac { n+1 -n}{\sqrt{n+1}+\sqrt{n}}=\\ \frac { 1}{\sqrt{n+1}+\sqrt{n}} \]

OpenStudy (anonymous):

Perfect. Was just googling conjugate as we spoke. Thanks very much for your time.

OpenStudy (anonymous):

yw